mathslover (mathslover):

IF : \[\LARGE{(\sqrt{2})^x+(\sqrt{3})^x=(\sqrt{13}^{\frac{x}{2}})}\] then the number of values of x = ?

OpenStudy (experimentx):

what's the ... on the left??

mathslover (mathslover):

what's the?? I didn't get you @experimentX

OpenStudy (experimentx):

(13)^(x/4) ??

mathslover (mathslover):

Yes ... \[\large{2^{\frac{x}{2}}+3^{\frac{x}{2}}=13^{\frac{x}{4}}}\]

OpenStudy (anonymous):

is this correct ? \[\large \sqrt{13}^{\frac{x}{2}}=13^{\frac{x}{4}}=(9+4)^{\frac{x}{4}} \ge 3^{\frac{x}{2}}+2^{\frac{x}{2}} \] and looking for the condition for which equality occurs...maybe using binomial theorem

mathslover (mathslover):

May be or may not be .. But the book is showing the answer as 1 ... can u explain more? why dont u check it out by checking the answer that you got with the book's ans.

mathslover (mathslover):

@mukushla any hint?

OpenStudy (anonymous):

im workin on it...

mathslover (mathslover):

take ur time then

OpenStudy (anonymous):

will this help?